# Sum frequency generation from real-time simulations in two-dimensional crystals

### v1

@@ Abstract
Sum frequency generation (SFG) and difference frequency generation (DFG) are second order nonlinear processes where two lasers with frequencies $\omega_{1}$ and $\omega_{2}$ combine to produce a response at frequency $\omega=\omega_{1}\pm\omega_{2}$. Compared with other nonlinear responses such as second-harmonic generation, SFG and DFG allow for tunability over a larger range. Moreover, the optical response can be enhanced by selecting the two laser frequencies in order to match specific electron-hole transitions.
Here, we propose a first-principles framework based on the real-time solution of an effective Schrödinger equation to calculate the SFG and DFG in various systems, such as bulk materials, 2D materials, and molecules. Within this framework, one can select from various levels of theory for the effective one-particle Hamiltonian to account for local-field effects and electron-hole interactions. To assess the approach, we calculate the SFG and DFG of two-dimensional crystals, h-BN and MoS2 monolayers, both within the independent-particle picture and including many-body effects. Additionally, we demonstrate that our approach can also extract higher-order response functions, such as field-induced second-harmonic generation. We provide an example using bilayer h-BN.
Sum frequency generation from real-time simulations in two-dimensional crystals
Mike N. Pionteck1*, Myrta Grüning2,3, Simone Sanna,1, Claudio Attaccalite3,4
@@ 1 Introduction
Sum frequency generation (SFG) and difference frequency generation (DFG) spectroscopy are powerful experimental techniques where the spectrum is the second-order nonlinear optical response $\chi^{(2)}$ resulting from the combination of two laser fields (see Fig. 1(a) and (b)). These techniques are highly sensitive to surfaces and interfaces [1, 2, 3]. In recent years, there has been a growing interest in the application of SFG/DFG in condensed matter physics. In particular, SFG/DFG was reported in layered MoS2 and related heterosystems using either band-filtered supercontinuum illumination [4, 5] or wavelength-dependent spectroscopy [6, 7, 5]. More interestingly, one can make the SFG dual resonant with the exciton, strongly enhancing its response function, as shown recently in two-dimensional (2D) materials [8]. Then, SFG can serve to explore exciton-exciton transitions as an alternative and complementary technique to pump and probe spectroscopy [9]. The design and interpretation of such experiments call for the development of theoretical approaches that can, for a specific material, capture both the nonlinear light-matter interaction and the many-body physics of excitons.
@@
So far, few theoretical studies have been reported in the literature on SFG and DFG in solids. The SFG was investigated using either simple two-band models [10] or from first-principles, using the Greenwood-Kubo formalism, within an independent-particle picture [8]—thus missing excitonic effects. In this work, we put forward a general approach to extract the SFG and DFG spectra from real-time simulations. We implement this approach within the first-principles framework of Ref. [11] in which the coupling of the electrons with the external electric field based on the Berry-phase formulation of the dynamical polarization [12] and many-body effects, including excitonic effects, are accounted for through an effective Hamiltonian [13]. We apply the approach to calculate the SFG and DFG of MoS2 and h-BN monolayers, both within the independent-particle picture and including excitonic effects. The approach presented is not limited to the SFG and DFG but allows an efficient calculation of other response functions, such as field-induced second-harmonic generation (FI-SHG) [14, 15]. FI-SHG involves applying an electric field, such as laser pulses, DC current, or an intense terahertz (THz) electric field, to a crystal and measuring the resulting second-harmonic intensity, which can provide important insights into the material properties. Centrosymmetric crystals, which have a null second harmonic response (in the dipole approximation), are of particular interest because the applied field breaks the symmetry and produces even-order harmonic radiation. For a static electric field a real-time approach to study FI-SHG has been proposed in Ref.[16]. Here we go a step further, and put forward a framework to simulate FI-SHG in presence of time-dependent pump fields. For a h-BN bilayer, which is a centrosymmetric crystal, we calculate the second harmonic response induced by a THz field breaking the inversion symmetry [Fig. 1(c)].
@@
The manuscript is organized as follows. In Sec. 2, we present the real-time approach used to obtain the dynamical polarization (Sec. 2.1) and provide a description of SFG, DFG, and FI-SHG through their Lehmann representations (Secs. 2.2-2.3). In Sec. 3, we detail and contrast the signal process techniques (discrete Fourier analysis and nonlinear least squares optimization) used to obtain the SFG/DFG and FI-SHG. After outlining the computational details in Sec. 4, we discuss in Sec. 5 the results obtained for the SFG and DFG of MoS2 and h-BN monolayers and for the FI-SHG in the h-BN bilayer.
@@ 2.1 Nonlinear response from real-time simulations
where $|v_{m\textbf{k}}\rangle=|v_{m\textbf{k}}(t)\rangle$ is the periodic part of the time-dependent Bloch states. On the right-hand side of Eq. (1), the second term describes the coupling with the external field in the dipole approximation. The coupling takes the form of a $\bf{k}$-derivative operator. The tilde indicates that the operator is gauge covariant, i.e. the solutions of Eq. (1) do not change under unitary rotation at $\bf{k}$ (see Ref. [12] for more details). $H^{\text{MB}}_{\mathbf{k}}$ is the effective many-body Hamiltonian. Different correlation effects can accounted for by constructing the corresponding effective Hamiltonian. The most accurate effective Hamiltonian we consider in this work is the time-dependent adiabatic GW (TD-aGW)[17] approximation which accounts for electron-hole interaction and local field effects:
$H^{\text{MB}}_{\mathbf{k}}\equiv H^{\text{KS}}_{\textbf{k}}+\Delta H_{\textbf{k}}+V_{h}(\mathbf{r})[\Delta\rho]+\Sigma_{\text{SEX}}[\Delta\gamma]\;,$ (2)
where $H^{\text{KS}}_{\textbf{k}}$ is the Hamiltonian of the unperturbed (zero-field) Kohn-Sham system[18], $\Delta H_{\textbf{k}}$ is the scissors operator applied to the Kohn-Sham eigenvalues, and the term $V_{h}(\mathbf{r})[\Delta\rho]$ is the real-time Hartree potential [11], responsible for the local field effects [19] arising from system inhomogeneities. The term $\Sigma_{\text{SEX}}$ is the screened-exchange self-energy that accounts for the electron-hole interaction [17]. $\Delta\rho$ and $\Delta\gamma$ are, respectively, the variations of the density and the density matrix induced by the external fields. In the limit of small perturbation Eqs. (1) and (2) reproduce the optical absorption calculated with the standard $GW$ + BSE approach,[19] as shown both analytically and numerically in Ref.[17]. The other approximation we consider is the independent particle approximation (IPA) that corresponds to $H^{\text{MB}}\equiv H^{\text{KS}}$.
@@
$P_{\parallel}(t)=-\frac{ef|\mathbf{a}|}{2\pi V}\text{Im log}\prod_{\textbf{k}}^{N_{\textbf{k}}-1}\ \text{det}S\left(\textbf{k},\textbf{k}+\mathbf{q};t\right),\;$ (3)
where $S(\textbf{k},\textbf{k}+\mathbf{q};t)$ is the overlap matrix between the time-dependent valence states $|v_{n\textbf{k}}\rangle$ and $|v_{m\textbf{k}+\textbf{q}}\rangle$, $V$ is the unit cell volume, $f$ is the spin degeneracy, $N_{\textbf{k}}$ is the number of $\bf{k}$-points along the polarization direction, and $\mathbf{q}=2\pi/(N_{\textbf{k}}{\mathbf{a}})$. Then, the $n$-order susceptibilities $\chi^{(n)}$ are extracted from the frequency-dependent polarization expanded in a power series of the incident fields as:
$P_{\alpha}(\omega)=\sum_{\beta=1}^{3}\sum_{i=1}^{2}\chi^{(1)}_{\alpha\beta}(\omega;\omega_{i})\mathcal{E}_{\beta}(\omega_{i})+\sum_{\beta,\gamma=1}^{3}\sum_{i,j=1}^{2}\chi^{(2)}_{\alpha\beta\gamma}(\omega;\omega_{i},\omega_{j})\mathcal{E}_{\beta}(\omega_{i})\mathcal{E}_{\gamma}(\omega_{j})+O(\mathcal{E}^{3})\;,$ (4)
@@ 2.2 Sum/Difference frequency generation
In Fig. 1(a),(b) we present a schematic representation of the SFG and DFG response functions studied in this work that correspond to $\chi^{(2)}(\omega,\omega_{i},\omega_{j})$ in Eq. (4) with $\omega=\omega_{i}\pm\omega_{j}$. In Sec. 3.2, we explain how the SFG and DFG are derived from the time-dependent polarization. Here, we discuss their Lehmann representation which is useful in the interpretation of the results.
@@
$\begin{split}\chi_{\alpha\beta\gamma}^{(2)}(\omega_{3};\omega_{1},\omega_{2})&\approx\frac{-ie^{3}}{m^{3}\tilde{\omega}_{3}\tilde{\omega}_{1}\tilde{\omega}_{2}}\sum_{\lambda\lambda^{\prime}}\Bigg{[}\frac{\alpha_{0\lambda}\beta_{\lambda\lambda^{\prime}}\gamma_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{3}-\Omega_{\lambda})}+\frac{\beta_{0\lambda}\gamma_{\lambda\lambda^{\prime}}\alpha_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}+\Omega_{\lambda})(\tilde{\omega}_{3}+\Omega_{\lambda^{\prime}})}\\ &-\frac{\gamma_{0\lambda}\alpha_{\lambda\lambda^{\prime}}\beta_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{2}+\Omega_{\lambda})}+\frac{\alpha_{0\lambda}\gamma_{\lambda\lambda^{\prime}}\beta_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{3}-\Omega_{\lambda})}\\ &+\frac{\gamma_{0\lambda}\beta_{\lambda\lambda^{\prime}}\alpha_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}+\Omega_{\lambda})(\tilde{\omega}_{3}+\Omega_{\lambda^{\prime}})}-\frac{\beta_{0\lambda}\alpha_{\lambda^{\prime}}\gamma_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{1}+\Omega_{\lambda})}\Bigg{]}\end{split}$ (5)
where $\Omega_{\lambda}$ are the excitation energies of the system, and $\alpha_{\lambda^{\prime}}$ refers to momentum matrix elements $\matrixelement{\lambda}{P_{\alpha}}{\lambda^{\prime}}$ between two excited states and similarly for $\beta_{\lambda\lambda^{\prime}}$ and $\gamma_{\lambda\lambda^{\prime}}$. Here, $\tilde{\omega}_{1}=\omega_{1}+i\eta$, $\tilde{\omega}_{2}=\omega_{2}+i\eta$ and $\tilde{\omega}_{3}=\tilde{\omega_{1}}+\tilde{\omega}_{2}$. $\mathbf{P}$ is the many-body momentum operator, i.e. $\mathbf{P}=\sum_{i}\mathbf{p}_{i}$ where $\mathbf{p}_{i}$ is the single-particle momentum operator acting on particle $i$ and $\eta$ is a small positive number that introduces dephasing/dissipation effects. An approximation of Eq. (5) can be derived by replacing many-body states and energies with excitonic ones [19]: $|\lambda\rangle\simeq|\Psi^{\text{exc}}_{\lambda}\rangle$ and $\Omega_{\lambda}\simeq E_{\lambda}$. When we consider the SFG case $\omega_{3}=\omega_{1}+\omega_{2}$, we retain only positive contributions to the $\chi^{(2)}$ and we get
@@
where $\lambda$ now indicates the excitonic state and $P_{\alpha,\lambda\lambda^{\prime}}=\langle\Psi^{\text{exc}}_{\lambda}|p_{\alpha}|\Psi^{\text{exc}}_{\lambda^{\prime}}\rangle$[9]. A similar procedure has been applied in the literature for the second-harmonic generation (SHG) case[21, 22]. From this formula we could expect strong peaks when $\omega_{1}+\omega_{2}$ is resonant with excitonic energy or when single laser frequencies $\omega_{1},\omega_{2}$ are resonant with an exciton. Note that the first and third terms have poles at both one-photon (e.g. $\omega_{1},\omega_{2}$) and two-photon ($\omega_{1}+\omega_{2}$) resonances. Finally, if $P_{\alpha,\lambda\lambda^{\prime}}$ is different from zero for $\lambda\neq\lambda^{\prime}$, there may also resonance with two distinct excitonic energies.

### v2

# A real-time approach to frequency-mixing spectroscopies: application to sum and difference frequency generation in two-dimensional crystals

@@ Abstract
We propose a computational framework to extract non-linear response functions from real-time simulations in the presence of more than one external field. We apply this approach to the calculation of sum frequency generation (SFG) and difference frequency generation (DFG). SFG and DFG are second-order nonlinear processes where two lasers with frequencies $\omega_{1}$ and $\omega_{2}$ combine to produce a response at frequency $\omega=\omega_{1}\pm\omega_{2}$. Compared with other nonlinear responses such as second-harmonic generation, SFG and DFG allow for tunability over a larger range. Moreover, the optical response can be enhanced by selecting the two laser frequencies in order to match specific electron-hole transitions.
To assess the approach, we calculate the SFG and DFG of two-dimensional crystals, h-BN and MoS2 monolayers, from real-time solution of an effective Schrödinger equation. Within the effective Schrödinger equation, one can select from various levels of theory for the effective one-particle Hamiltonian to account for local-field effects and electron-hole interactions. We compare results obtained within the independent-particle picture and including many-body effects. Such comparison allows us to identify and characterize excitonic features in the obtained spectra. Additionally, we demonstrate that our approach can also extract higher-order response functions, such as field-induced second-harmonic generation. We provide an example using the h-BN bilayer.
A real-time approach to frequency-mixing spectroscopies: application to sum and difference frequency generation in two-dimensional crystals
Mike N. Pionteck1*, Myrta Grüning2,3, Simone Sanna1, Claudio Attaccalite3,4
@@ 1 Introduction
Sum frequency generation (SFG) and difference frequency generation (DFG) spectroscopy are powerful experimental techniques where the spectrum is the second-order nonlinear optical response $\chi^{(2)}$ resulting from the combination of two laser fields (see Fig. 1(a) and (b)). These techniques are highly sensitive to surfaces and interfaces [1, 2, 3]. In recent years, there has been a growing interest in the application of SFG/DFG in condensed matter physics. In particular, SFG/DFG was reported in layered MoS2 and related heterosystems using either band-filtered supercontinuum illumination [4, 5] or wavelength-dependent spectroscopy [6, 7, 5, 8]. More interestingly, one can make the SFG dual resonant with the exciton, strongly enhancing its response function, as shown recently in two-dimensional (2D) materials [9]. Then, SFG can serve to explore exciton-exciton transitions as an alternative and complementary technique to pump and probe spectroscopy [10]. The design and interpretation of such experiments call for the development of theoretical approaches that can, for a specific material, capture both the nonlinear light-matter interaction and the many-body physics of excitons.
@@
So far, few theoretical studies have been reported in the literature on SFG and DFG in solids. The SFG was investigated using either simple two-band models [11] or from first-principles, using the Greenwood-Kubo formalism, within an independent-particle picture [9]—thus missing excitonic effects. In this work, we put forward a general approach to extract the SFG and DFG spectra from real-time simulations. We implement this approach within the first-principles framework of Ref. [12] in which the coupling of the electrons with the external electric field based on the Berry-phase formulation of the dynamical polarization [13] and many-body effects, including excitonic effects, are accounted for through an effective Hamiltonian [14]. We apply the approach to calculate the SFG and DFG of MoS2 and h-BN monolayers, both within the independent-particle picture and including excitonic effects. The approach presented is not limited to the SFG and DFG but allows an efficient calculation of other response functions, such as field-induced second-harmonic generation (FI-SHG) [15, 16]. FI-SHG involves applying an electric field, such as laser pulses, DC current, or an intense terahertz (THz) electric field, to a crystal and measuring the resulting second-harmonic intensity, which can provide important insights into the material properties. Centrosymmetric crystals, which have a null second-harmonic response (in the dipole approximation), are of particular interest because the applied field breaks the symmetry and produces even-order harmonic radiation. For a static electric field a real-time approach to study FI-SHG has been proposed in Ref.[17]. Here we go a step further, and put forward a framework to simulate FI-SHG in the presence of time-dependent pump fields. For a h-BN bilayer, which is a centrosymmetric crystal, we calculate the second-harmonic response induced by a THz field breaking the inversion symmetry [Fig. 1(c)].
@@
The manuscript is organized as follows. In Sec. 2, we present the real-time approach used to obtain the dynamical polarization (Sec. 2.1) and provide a description of SFG, DFG, and FI-SHG through their Lehmann representations (Secs. 2.2-2.3). In Sec. 3, we detail and contrast the signal process techniques (discrete Fourier analysis and least squares optimization) used to obtain the SFG/DFG and FI-SHG. After outlining the computational details in Sec. 4, we discuss in Sec. 5 the results obtained for the SFG and DFG of MoS2 and h-BN monolayers and for the FI-SHG in the h-BN bilayer.
@@ 2.1 Nonlinear response from real-time simulations
where $|v_{m\textbf{k}}\rangle=|v_{m\textbf{k}}(t)\rangle$ is the periodic part of the time-dependent Bloch states. On the right-hand side of Eq. (1), the second term describes the coupling with the external field in the dipole approximation. The coupling takes the form of a $\bf{k}$-derivative operator. The tilde indicates that the operator is gauge covariant, i.e. the solutions of Eq. (1) do not change under unitary rotation at $\bf{k}$ (see Ref. [13] for more details). $H^{\text{MB}}_{\mathbf{k}}$ is the effective many-body Hamiltonian. Different correlation effects can be accounted for by constructing the corresponding effective Hamiltonian. The simplest approximation for the effective Hamiltonian we consider in this work is the independent particle approximation (IPA) describing the unperturbed (zero-field) Kohn-Sham system[18]
$H^{\text{MB}}\equiv H^{\text{KS}}=-\frac{\hbar^{2}}{2m}\sum_{i}\nabla_{i}^{2}+V_{\mathrm{eI}}+V_{\mathrm{h}}[\rho^{0}]+V_{\mathrm{xc}}[\rho^{0}]=\sum_{i}\varepsilon_{n,\mathbf{k}}\left|u_{n,\mathbf{k}}\right\rangle\left\langle u_{n,\mathbf{k}}\right|,$ (2)
where $\hat{V}_{\mathrm{eI}}$ is the electron-ion interaction, $\hat{V}_{\mathrm{H}}[\rho^{0}]$ and $\hat{V}_{\mathrm{xc}}[\rho^{0}]$ are respectively the Hartree and exchange-correlation potentials evaluated at the unperturbed ground-state electron density $\rho^{0}$, $\varepsilon_{n,\mathbf{k}}$ the Kohn-Sham energies (eigenvalues) and $|u_{n,\mathbf{k}}\rangle$ the periodic part of the Kohn-Sham wavefunctions. The Hamiltonian in Eq. (2) has two main shortcomings. First, the band gaps obtained from the Kohn-Sham energies are systematically underestimated—yielding to an overall underestimation of the optical gap. Second, the response of the density-functional potentials to the change in the density is missing and so are local-field effects and quasi-particle excitations. The former shortcoming is corrected by introducing a state-dependent scissor operator term,
$\Delta\hat{H}_{\mathbf{k}}=\sum_{n}\Delta_{n\mathbf{k}}\left|u_{n,\mathbf{k}}\right\rangle\left\langle u_{n,\mathbf{k}}\right|,$ (3)
where state-dependent $\Delta_{n\mathbf{k}}$ is the scissor operator. That is, considering the Kohn-Sham energies as the zero-order approximation to quasiparticle energies, the scissor operator introduces a first order correction. Such corrections can be obtained from first-principles within the so-called $GW$ approximation derived from many-body perturbation theory[19] (a simpler approach is to have a single scissor operator chosen to reproduce e.g. the experimental fundamental band gap). To address the second shortcoming, we include the fluctuation of the Hartree potential (Eq. (2)) generated by the variation of the electron density $\Delta\rho=\rho-\rho^{0}$ induced by the external fields [12], and the screened-exchange self-energy term $\Sigma_{\text{SEX}}[\Delta\gamma]$ where $\Delta\gamma$ is the fluctuation of the density matrix [20]. This latter term is a nonlocal operator which accounts for the screened electron-hole attraction and introduces excitonic effects. We so obtain the following effective Hamiltonian,
$H^{\text{MB}}_{\mathbf{k}}\equiv H^{\text{KS}}_{\textbf{k}}+\Delta H_{\textbf{k}}+V_{\mathrm{h}}(\mathbf{r})[\Delta\rho]+\Sigma_{\text{SEX}}[\Delta\gamma]\;,$ (4)
which is the most accurate Hamiltonian considered in this work and is known as time-dependent adiabatic GW (TD-aGW) [^1] Notice that the TD-aGW approximation was called TD-BSE, TD-SEX or TD-HSEX in previous publications [10, 20, 21]. approximation [20]. This Hamiltonian (Eq. (4)) is built in such a way to reproduce the band structure and the response functions of the the standard $GW$ + BSE approach [22, 19]. This has been shown in the limit of small perturbation both analytically and numerically in Ref. [20].
@@
$\mathcal{P}_{\parallel}(t)=-\frac{ef|\mathbf{a}|}{2\pi V}\text{Im log}\prod_{\textbf{k}}^{N_{\textbf{k}}-1}\ \text{det}S\left(\textbf{k},\textbf{k}+\mathbf{q};t\right),\;$ (5)
where $S(\textbf{k},\textbf{k}+\mathbf{q};t)$ is the overlap matrix between the time-dependent valence states $|v_{n\textbf{k}}\rangle$ and $|v_{m\textbf{k}+\textbf{q}}\rangle$, $V$ is the unit cell volume, $f$ is the spin degeneracy, $N_{\textbf{k}}$ is the number of $\bf{k}$-points along the polarization direction, and $\mathbf{q}=2\pi/(N_{\textbf{k}}{\mathbf{a}})$. Then, the $n$-order susceptibilities $\chi^{(n)}$ are extracted from the frequency-dependent polarization expanded in a power series of the two incident fields as:
$\begin{split}\mathcal{P}_{\alpha}(\omega)=\sum_{\beta=1}^{3}\sum_{i=1}^{2}\chi^{(1)}_{\alpha\beta}(\omega;\omega_{i})\mathcal{E}_{\beta}(\omega_{i})+\sum_{\beta,\gamma=1}^{3}\sum_{i,j=1}^{2}\chi^{(2)}_{\alpha\beta\gamma}(\omega;\omega_{i},\omega_{j})\mathcal{E}_{\beta}(\omega_{i})\mathcal{E}_{\gamma}(\omega_{j})\\ +\sum_{\beta,\gamma,\delta=1}^{3}\sum_{i,j,k=1}^{2}\chi^{(3)}_{\alpha\beta\gamma\delta}(\omega;\omega_{i},\omega_{j},\omega_{j})\mathcal{E}_{\beta}(\omega_{i})\mathcal{E}_{\gamma}(\omega_{j})\mathcal{E}_{\delta}(\omega_{k})+O(\mathcal{E}^{4})\;,\end{split}$ (6)
@@ 2.2 Sum/Difference frequency generation
In Fig. 1(a),(b) we present a schematic representation of the SFG and DFG response functions studied in this work which correspond to $\chi^{(2)}(\omega,\omega_{i},\omega_{j})$ in Eq. (6) with $\omega=\omega_{i}\pm\omega_{j}$. In Sec. 3.2, we explain how the SFG and DFG are derived from the time-dependent polarization. Here, we discuss their Lehmann representation which is useful in the interpretation of the results.
@@
$\begin{split}\chi_{\alpha\beta\gamma}^{(2)}(\omega_{3};\omega_{1},\omega_{2})&=\frac{-ie^{3}}{m^{3}\tilde{\omega}_{3}\tilde{\omega}_{1}\tilde{\omega}_{2}}\sum_{\lambda\lambda^{\prime}}\Bigg[\frac{\alpha_{0\lambda}\beta_{\lambda\lambda^{\prime}}\gamma_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{3}-\Omega_{\lambda})}+\frac{\beta_{0\lambda}\gamma_{\lambda\lambda^{\prime}}\alpha_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}+\Omega_{\lambda})(\tilde{\omega}_{3}+\Omega_{\lambda^{\prime}})}\\ &-\frac{\gamma_{0\lambda}\alpha_{\lambda\lambda^{\prime}}\beta_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{2}+\Omega_{\lambda})}+\frac{\alpha_{0\lambda}\gamma_{\lambda\lambda^{\prime}}\beta_{\lambda^{\prime}0}}{(\tilde{\omega}_{1}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{3}-\Omega_{\lambda})}\\ &+\frac{\gamma_{0\lambda}\beta_{\lambda\lambda^{\prime}}\alpha_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}+\Omega_{\lambda})(\tilde{\omega}_{3}+\Omega_{\lambda^{\prime}})}-\frac{\beta_{0\lambda}\alpha_{\lambda^{\prime}}\gamma_{\lambda^{\prime}0}}{(\tilde{\omega}_{2}-\Omega_{\lambda^{\prime}})(\tilde{\omega}_{1}+\Omega_{\lambda})}\Bigg]\end{split}$ (7)
[…]
@@
where $\lambda$ now indicates the excitonic state and $P_{\alpha,\lambda\lambda^{\prime}}=\langle\Psi^{\text{exc}}_{\lambda}|p_{\alpha}|\Psi^{\text{exc}}_{\lambda^{\prime}}\rangle$[10]. A similar procedure has been applied in the literature for the second-harmonic generation (SHG) case[24, 25]. From this formula we could expect strong peaks when $\omega_{1}+\omega_{2}$ is resonant with the excitonic energy or when single laser frequencies $\omega_{1},\omega_{2}$ are resonant with an exciton. Note that the first and third terms have poles at both one-photon (e.g. $\omega_{1},\omega_{2}$) and two-photon ($\omega_{1}+\omega_{2}$) resonances. Finally, if $P_{\alpha,\lambda\lambda^{\prime}}$ is different from zero for $\lambda\neq\lambda^{\prime}$, there may also be resonances with two distinct excitonic energies.